mother weight is 150lbs and she is 5ft 4inches . the nurse obtains measurements of 65kg how many lbs did the mother lose
Answer #1

Weight of the mother = 150lbs

that implies, weight of the mother (in kgs) = (150* 0.453)kg

                                                                    = 68.03 kgs                                         (.: 1lb= 0.453 kg)

Loss of weight = (68.03 - 65)

                         = 3.03 kgs

                         = 3.03 * 2.2                                                                                     (.: 1kg = 2.20lbs)

                        = 6.68 lbs.

So, the mother lost weight of 6.68 lbs.